If a consumer is consuming two commodities X and Y and his utility function U (X,Y)=2XY+6.if the price of the two commodity are 2 and 6 respectively, and the consumer has a total income of 80 birr to be spent on the two goods, a)find the utility maximizing Quantity of good X and Y. b) find the MRSxy at equilibrium. c) find the MRSyx at equilibrium.


1
Expert's answer
2022-02-15T15:59:03-0500

"U(X,Y)= 2XY+6"

"P_x= 2"

"P_y=6"

I= 80

Budget line

2x+ 6Y= 80

a) Utility maximization Quantity

"\frac{Mu_x}{Mu_y}=\frac{P_x}{P_y}"

"\frac{2Y}{2X}=\frac{2}{6}"


12Y= 4X

Y= "\frac{1}{3}"X

X= 3Y

Plug in the above in the budget line

2(3Y)+ 6Y= 80

12Y= 80

Y*= 6.67

2X+ "6(\frac{1}{3}X)"= 80

4X= 80

X*= 20

b) MRS"_{xy}= \frac{Mu_x}{Mu_y}"

"= \frac{2Y}{2X}= \frac{2(\frac{20}{3})}{2(20)}"

="\frac{40}{3}\times\frac{1}{40}= 0.33"


c) "MRS_{yx}= \frac{Mu_y}{Mu_x}= \frac{2X}{2Y}"

= "\frac{20\times 20}{2\times \frac{20}{3}}" = 3